A car is travelling at a constant speed of 45 m/s passes a topper behind a billboard. one second of the car passes the billboard the trooper sets out from the billboard to catch it,accelerating at a constant speed of 3 m/s. how long does it take her to overtake the car?
1
Expert's answer
2020-01-31T10:44:57-0500
"s=v(t+1)=0.5at^2"

"45(t+1)=0.5(3)t^2"

"t=31\ s"